### v1

@@ Abstract
Figure 2: Statistics of our PangeaIns, which comprises 6M multimodal instructions in 39 languages. PangeaIns includes general instructions, document and chart question answering, captioning, domain-specific, culturally relevant, and text-only instructions. To address the data scarcity of multilingual multimodal instructions, we employ three strategies: 1) translating high-quality English instructions into multiple languages, filtering culturally relevant images; 2) generating culturally aware instructions, and 3) collecting and re-purposing existing open-source multimodal datasets.
@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; Deitke et al., 2024; OpenAI, 2024b; Team et al., 2023) have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and complex reasoning (Yue et al., 2024a; b). Despite this rapid progress in their reasoning capabilities, a critical flaw persists: the overwhelming focus on English- and western-centric training datasets and evaluation benchmarks (Liu et al., 2021; Song et al., 2023).
This homogeneity results in a lack of representation for the vast majority of the world’s languages and diverse cultural contexts (Yu et al., 2022). Consequently, models predominately trained on such data exhibit: (a) diminished performance in multilingual settings (Blasi et al., 2022) with poor tokenization further leading to higher inference costs (Ahia et al., 2023); (b) generate outputs misaligned with the socio-cultural norms of underrepresented languages (AlKhamissi et al., 2024); and (c) lack the ability to recognize objects from geographically diverse regions (Ramaswamy et al., 2024) or rare objects belonging to the long-tail (Gupta et al., 2019). With the increased adoption of these models into real-world applications across the globe, there’s an urgent need to develop multilingual MLLMs that equitably serve a diverse set of users. Few efforts have been made to develop multilingual MLLMs (Geigle et al., 2024; Rasheed et al., 2025), however, their performance still exhibits inequalities across languages and lacks evaluation of cultural understanding.
@@
Our results demonstrate Pangea abilities in both English and multilingual scenarios, significantly outperforming existing open-source MLLMs across PangeaBench datasets, surpassing the best open MLLMs by 7.3 points on English tasks and 10.8 points on multilingual tasks on average. Notably, Pangea excels in multilingual and multi-cultural understanding, as evidenced by its performance on xChat and CVQA benchmarks. Pangea also matches or outperforms state-of-the art proprietary LLMs, namely Gemini-1.5-Pro and GPT4o, on several tasks. However, some performance gaps remain in multimodal chat and complex reasoning, shedding light on the need for further improvements in open multimodal LLMs. We discuss key insights from Pangea’s training, including the scaling effect of instructions, the role of English data, and the impact of language-specific training proportions on performance. Additionally, we explore preliminary methods to improve multilingual OCR. We fully open-source PangeaIns, PangeaBench, Pangea models, and code, enabling advances in culturally inclusive MLLMs across diverse languages.
@@ 2.1 Machine Translated Instructions
Overall, we found the Gemini 1.5 Pro translation to be satisfactory—providing a fast, cost-effective alternative to human annotation, especially for scaling across multiple languages. However, we acknowledge that machine translation still has some limitations, particularly in handling nuanced contexts and cultural subtleties.
@@ 2.2 Multicultural Understanding Instructions
While machine translation enables us to scale across multiple languages, data translated from English is still Anglo-centric in its coverage of cultural concepts (Yu et al., 2022). To address this, we developed a pipeline focused on creating instructions for multicultural understanding. Both visual and textual elements can convey deep cultural significance, and our goal is to design a dataset that allows models to not only recognize these nuances but also respond appropriately across various cultural contexts. The pipeline of creating multicultural understanding instructions is shown in Figure 3.
@@
  • Avoiding Overrepresentation: To maintain a balanced representation, we downsampled images from frequently occurring subjects, such as objects, materials, and clothing, to avoid skewing the dataset toward specific topics or regions. Additionally, an accessibility check was conducted, which removed around 30% of the remaining samples due to image download or other issues. Ultimately, we curated a final set of 1 million high-quality, culturally specific images that form the foundation of our dataset.
@@ 2.4 Dataset Statistics
Language and Task Distribution: PangeaIns features an extensive and balanced distribution of languages, tasks, and cultural contexts (as shown in Figure 2). We empirically keep the final language ratio of English to Multilingual as 40%:60% as we found a significant portion of English data plays an important role in cross-lingual transfer. See more discussions about the ratio in section 5 and Figure 6. The inclusion of diverse multimodal instructions ensures that the model develops a deeper understanding of varied linguistic and cultural environments. Examples of training samples from different languages and categories are provided in Appendix G.
@@
Category Tasks Datasets Forms Size Languages
Multimodal Multimodal Chat xChatBench Long 400 zh,en,hi,id,ja,rw,ko,es
M-LlavaBench Long 600 ar,bn,zh,fr,hi,ja,ru,es,ur,en
Captioning XM100 Long 3.6K 36 languages
Cultural Understanding CVQA MC 21K en,zh,ko,mn,ja,id,jv,min,su,
MaRVL Short 6K id,sw,ta,tr,zh
Multilingual VQA xGQA Short 77K en,de,pt,ru,id,bn,ko,zh
MaXM MC 2K hi,th,zh,fr,en,iw,ro
Reasoning (Multi-subject) xMMMU Short/MC 3K en,ar,fr,hi,id,ja,pt
M3Exam MC 3K en,zh,it,pt,vi,th,af
Text-only QA TyDiQA Short 5.1K ar,ru,bn,te,fi,sw,ko,id,en
Translation FLORES-Sub Long 18K ar,en,fr,de,hi,id,iw,ja,pt,ro,tr
Reasoning (Multi-subject, Commonsense, Math) MMMLU MC 197K ar,bn,de,es,fr,hi,id,it,ja,ko,pt,sw,yo,zh
XStoryCloze MC 21K en,ar,es,eu,hi,id,my,ru,sw,te,zh
MGSM Open 3K bn,de,en,es,fr,ja,ru,sw,te,th,zh
@@ 3.1 Overview of PangeaBench
To assess the capabilities of Pangea across a variety of languages, cultures, and task types, we have developed PangeaBench, a comprehensive multilingual and multimodal evaluation suite. PangeaBench integrates diverse benchmarks that encompass both multimodal and text-only tasks, enabling a holistic evaluation of Pangea’s performance in cross-lingual, cross-cultural, and multimodal contexts. Each task within PangeaBench is designed to probe specific aspects of Pangea’s proficiency, ensuring robust testing across a wide range of scenarios.
@@ 3.2 Multimodal Tasks
The multimodal tasks in PangeaBench are categorized as follows: Multimodal Chat, Captioning, Cultural Understanding, Multilingual Visual Question Answering (VQA), and Multi-Subject Reasoning. These tasks incorporate datasets from PangeaBench to ensure comprehensive testing of Pangea’s multimodal capabilities. The overview and examples of PangeaBench from each task are shown in Figure 4.
Multimodal Chat. This task tests the model’s ability to engage in natural, dynamic real-world conversations involving both text and images. In this category, Multilingual LlavaBench (Rasheed et al., 2025) (M-LlavaBench for short) stands as the only benchmark for evaluating multilingual long-form generation capabilities from MLLMs. Moreover, following the evaluation pipeline from Zheng et al. (2023) and Liu et al. (2023b), M-LlavaBench uses a coarse-grained evaluation criteria (i.e., “Please rate the helpfulness, relevance, accuracy, level of details of their responses.”). Prior works have shown that employing such coarse-grained evaluation criteria might lead to automatic evaluation results that diverge from how humans would evaluate them (Ye et al., 2023; Kim et al., 2023; Lee et al., 2024; Kim et al., 2024a; b). To assess our baselines based on a more accurate evaluation pipeline with fine-grained evaluation criteria on diverse scenarios, we additionally annotate a new multilingual multimodal generation benchmark called the xChatBench. An additional explanation of the annotation process of xChatBench is included in Appendix E. For the multimodal chat scenarios, we found that many English-centric models tend to respond in English regardless of the query language. This behavior is problematic, as it undermines the fundamental capability of a multilingual model, which should ideally respond in the language of the query. To address this, we implemented a strict evaluation criterion where such responses were penalized and assigned a score of 0. We believe this is crucial, as users may not understand English, and failing to respond in the appropriate language can hinder effective communication and user experience.
@@ 3.3 Text-Only Multilingual Datasets
$\Delta$ over SoTA Open +0.4 +10.8 -3.5 +7.1 -11.7 +18.3 -5.8 -4.2 +11.6 +12.3
@@
$\Delta$ over Best Open Model -0.2 +3.9 -0.1 +12.0 0.0 +9.3 -0.8 +2.3 +1.0 -3.7
@@
4 Experiments
4.1 Experimental Setup
We train Pangea on PangeaIns, our multilingual multimodal dataset comprising 6 million samples across 39 languages. The model is based on LLaVA-Next architecture (Liu et al., 2024) with Qwen2-7B-Instruct (Yang et al., 2024) as the language model backbone. We employ a learning rate of 2e-5, a batch size of 512, coupled with a cosine decay schedule with 0.03 warmup steps. We train the model for 1 epoch.
For evaluation, we compare Pangea against several state-of-the-art open source baselines, including English-centric models Llava-1.5-7B (Liu et al., 2023a), Llava-Next-7B (Liu et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), Cambrian-8B (Tong et al., 2024), Llava-OV-7B (Li et al., 2024b), Molmo-7B-D (Deitke et al., 2024) Llama3.2-11B (Dubey et al., 2024) and multilingual models PaliGemma-3B (Beyer et al., 2024), PALO-7B (Rasheed et al., 2025), mBLIP mT0-XL and mBLIP BLOOMZ (Geigle et al., 2024). We also consider two text-only LLMs baselines Vicuna-1.5-7B (Zheng et al., 2023) and Qwen2-7B-Instruct (Yang et al., 2024), which are the backbones of Llava-Next and our Pangea respectively. We integrate our multimodal tasks in PangeaBench into lmms-eval (Li et al., 2024a), [^1] https://github.com/EvolvingLMMs-Lab/lmms-eval a multimodal evaluation package that supports many English multimodal benchmarks. We use lm-evaluation-harness (Biderman et al., 2024) [^2] https://github.com/EleutherAI/lm-evaluation-harness to evaluate text-only tasks. We follow the original paper for their best models’ prompts in different tasks, and mostly reproduce their original numbers on datasets reported in the original papers.
@@ 5 Discussion
Role of English Data. In multilingual scenarios, English data plays a pivotal role in cross-lingual transfer. To investigate this, we sampled 500K examples from the translated data described in subsection 2.1, ensuring a consistent data distribution. We varied the ratio of English data while keeping the total number of training samples fixed at 500K. For the 17 multilingual languages in the translated subset, we evenly distributed the number of samples across languages. As shown in Figure 6 , English performance generally improves as the percentage of English data increases. More surprisingly, using no English data (full multilingual data) results in relatively lower multilingual performance. As we introduce more English data, multilingual performance improves, peaking at 38.7% with 40% English. However, performance drops sharply when English data reaches 100%. This suggests that English data aids cross-lingual transfer, however, over-reliance on it harms multilingual performance.
@@
Multimodal Chat in the Wild. One important application of MLLMs is to answer users’ queries in the wild. Here, we show the outputs of Pangea for the multimodal chat queries from our xChatBench. As shown in Appendix Figure 10, 11, 12, 13, 14, 15, Pangea successfully interprets the figures in different tasks and generates fluent and readable in certain languages. These qualitative examples further demonstrate the remarkable visual understanding ability of Pangea in multilingual contexts. On the other hand, we also identified a few bad cases shown in Figure 16, 17. Despite generating relevant responses to the queries, the model does not capture the key details of the images due to the lack of knowledge, which points out potential improvement directions in the future.
Figure 8: A preliminary exploration of multilingual OCR.
Preliminary Explorations of Multilingual OCR. Multilingual OCR emerged as a particularly challenging aspect of Pangea’s functionality. We made efforts to enhance its multilingual OCR capabilities. Specifically, we constructed a dataset of 500K multilingual OCR instructions spanning 10 languages, with 50K examples per language, sourced from web user interfaces. Webpages naturally serve as image-rich environments containing abundant text, and by capturing screenshots of websites from various countries in different languages, we were able to gather a substantial number of OCR images. Further details on the construction of the multilingual OCR training dataset are available in Appendix I.

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; 2024; Dubey et al., 2024; Deitke et al., 2024; OpenAI, 2024b; Team et al., 2023) have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and complex reasoning (Yue et al., 2024a; b). Despite this rapid progress in their reasoning capabilities, a critical flaw persists: the overwhelming focus on English- and western-centric training datasets and evaluation benchmarks (Liu et al., 2021; Song et al., 2023).
This homogeneity results in a lack of representation for the vast majority of the world’s languages and diverse cultural contexts (Yu et al., 2022). Consequently, models predominately trained on such data exhibit: (a) diminished performance in multilingual settings (Blasi et al., 2022) with poor tokenization further leading to higher inference costs (Ahia et al., 2023); (b) generate outputs misaligned with the socio-cultural norms of underrepresented languages (AlKhamissi et al., 2024); and (c) lack the ability to recognize objects from geographically diverse regions (Ramaswamy et al., 2024) or rare objects belonging to the long-tail (Gupta et al., 2019). With the increased adoption of these models into real-world applications across the globe, there’s an urgent need to develop multilingual MLLMs that equitably serve a diverse set of users. Few efforts have been made to develop multilingual MLLMs (Geigle et al., 2024b; Rasheed et al., 2025), however, their performance still exhibits inequalities across languages and lacks evaluation of cultural understanding.
@@
Our results demonstrate Pangea’s abilities in both English and multilingual scenarios, significantly outperforming existing open-source MLLMs across PangeaBench datasets, surpassing the best open MLLMs by 0.4 points on English tasks and 10.9 points on multilingual tasks on average. Notably, Pangea excels in multilingual and multi-cultural understanding, as evidenced by its performance on xChat, CVQA, and MaRVL benchmarks. Pangea also matches or outperforms state-of-the art proprietary LLMs, namely Gemini-1.5-Pro and GPT4o, on several tasks such as XGQA. However, some performance gaps remain in multimodal chat and complex reasoning, shedding light on the need for further improvements in open multimodal LLMs. We discuss key insights from Pangea’s training, including the scaling effect of instructions, the role of English data, and the impact of language-specific training proportions on performance. Additionally, we explore preliminary methods to improve multilingual OCR. We fully open-source PangeaIns, PangeaBench, Pangea models, and code, to advance culturally inclusive MLLMs across diverse languages.
Figure 2: Statistics of our PangeaIns, which comprises 6M multimodal instructions in 39 languages. The distribution of multilingual training samples shows the percentage of instances for each language among the multilingual instructions. PangeaIns includes general instructions, document and chart question answering, captioning, domain-specific, culturally relevant, and text-only instructions. To address the data scarcity of multilingual multimodal instructions, we employ three strategies: 1) translating high-quality English instructions into multiple languages, filtering culturally relevant images; 2) generating culturally aware instructions, and 3) collecting and re-purposing existing open-source multimodal datasets.
@@ 2.1 Machine Translated Instructions
Overall, Gemini 1.5 Pro’s translation seems satisfactory, providing a fast, cost-effective alternative to human annotation, especially for scaling across languages. However, we acknowledge that machine translation still has limitations, particularly in handling nuanced contexts and cultural subtleties.
@@ 2.2 Multicultural Understanding Instructions
While machine translation enables scaling across multiple languages, data translated from English is still Anglo-centric in coverage of cultural concepts (Yu et al., 2022). To address this, we developed a pipeline focused on creating instructions for multicultural understanding. Both visual and textual elements can convey deep cultural significance, and our goal is to design a dataset that allows models to not only recognize these nuances but also respond appropriately across various cultural contexts. The pipeline of creating multicultural understanding instructions is shown in Figure 3.
@@
  • Avoiding Overrepresentation: To maintain a balanced representation, we downsampled images from frequently occurring subjects, such as objects, materials, and clothing, to avoid skewing the dataset toward specific topics or regions. Then, we conducted an accessibility check, removing 30% of the remaining samples due to image download or other issues. Ultimately, we curated a final set of 1M high-quality, culturally specific images, forming the foundation of PangeaIns.
@@ 2.4 Dataset Statistics
Language and Task Distribution: PangeaIns features an extensive and balanced distribution of languages, tasks, and cultural contexts (as shown in Figure 2). We empirically keep the final language ratio of English to Multilingual as 40%:60% as we found a significant portion of English data plays an important role in cross-lingual transfer. See more discussions about the ratio in section 5 and Figure 6. The inclusion of diverse multimodal instructions ensures that the model develops a deeper understanding of varied linguistic and cultural environments. Examples of training samples from different languages and categories are provided in Appendix H.
@@
Category Tasks Datasets Forms Size Languages Metric
Multimodal Multimodal Chat xChatBench Long 400 zh,en,hi,id,ja,rw,ko,es LLM-as-Judge
M-LlavaBench Long 600 ar,bn,zh,fr,hi,ja,ru,es,ur,en LLM-as-Judge
Captioning XM100 Long 3.6K 36 languages ROUGE-L
Cultural Understanding CVQA MC 21K en,zh,ko,mn,ja,id,jv,min,su Accuracy
MaRVL Short 6K id,sw,ta,tr,zh Accuracy
Multilingual VQA xGQA Short 77K en,de,pt,ru,id,bn,ko,zh Accuracy
MaXM MC 2K hi,th,zh,fr,en,iw,ro Accuracy
Reasoning (Multi-subject) xMMMU Short/MC 3K en,ar,fr,hi,id,ja,pt Accuracy
M3Exam MC 3K en,zh,it,pt,vi,th,af Accuracy
Text-only QA TyDiQA Short 5.1K ar,ru,bn,te,fi,sw,ko,id,en Accuracy
Translation FLORES-Sub Long 18K ar,en,fr,de,hi,id,iw,ja,pt,ro,tr ChrF
Reasoning (Multi-subject, Commonsense, Math) MMMLU MC 197K ar,bn,de,es,fr,hi,id,it,ja,ko,pt,sw,yo,zh Accuracy
XStoryCloze MC 21K en,ar,es,eu,hi,id,my,ru,sw,te,zh Accuracy
MGSM Open 3K bn,de,en,es,fr,ja,ru,sw,te,th,zh Accuracy
@@ 3.1 Overview of PangeaBench
To assess the capabilities of Pangea across a variety of languages, cultures, and task types, we have developed PangeaBench, a comprehensive multilingual and multimodal evaluation suite. PangeaBench integrates diverse benchmarks that encompass both multimodal and text-only tasks, enabling a holistic evaluation of Pangea’s performance in cross-lingual, cross-cultural, and multimodal contexts. Each task within PangeaBench is designed to probe specific aspects of Pangea’s proficiency, ensuring robust testing across a wide range of scenarios. All tasks in PangeaBench are evaluated under a zero-shot setting.
@@ 3.2 Multimodal Tasks
The multimodal tasks in PangeaBench are categorized as follows: Multimodal Chat, Captioning, Cultural Understanding, Multilingual Visual Question Answering (VQA), and Multi-Subject Reasoning. We incorporate these tasks in PangeaBench to ensure comprehensive testing of Pangea’s multimodal capabilities. The overview and examples of PangeaBench from each task are shown in Figure 4.
Multimodal Chat. This task tests the model’s ability to engage in natural and dynamic real-world conversations involving both text and images. In this category, Multilingual LlavaBench (Rasheed et al., 2025) (M-LlavaBench for short) stands as the only benchmark for evaluating multilingual long-form generation capabilities from MLLMs. Moreover, following the evaluation pipeline from Zheng et al. (2023) and Liu et al. (2023b), M-LlavaBench uses a coarse-grained evaluation criteria (for example, “Please rate the helpfulness, relevance, accuracy, level of details of their responses.”). Previous works have shown that employing such coarse-grained evaluation criteria might lead to automatic evaluation results that diverge from how humans would evaluate them (Ye et al., 2023; Kim et al., 2023; Lee et al., 2024; Kim et al., 2024a; b). To assess our baselines based on a more accurate evaluation pipeline with fine-grained evaluation criteria on diverse scenarios, we additionally annotate a new multilingual multimodal generation benchmark called the xChatBench, included in the multimodal chat category of PangeaBench. A more detailed explanation of the annotation process of xChatBench is included in Appendix E.
@@ 3.3 Text-Only Multilingual Datasets
4 Experiments
4.1 Experimental Setup
Stages Pretraining Finetuning
Training Data
Dataset LLaVA LCS-558K PangeaIns
#Samples 558K 6M
Model
Trainable Projector (20M) Full Model (8B)
Training
Batch Size 128 128
LR: $\psi_{\text{vision}}$ 1 $\times$ $10^{-3}$ 2 $\times$ $10^{-6}$
LR: $\{\theta_{\text{proj}},\phi_{\text{LLM}}\}$ 1 $\times$ $10^{-3}$ 2 $\times$ $10^{-5}$
Epoch 1 1
GPU Hours (H100) 32 1344
Table 1: Pangea’s training configurations.
We train Pangea on PangeaIns, our multilingual multimodal dataset comprising 6 million samples across 39 languages. The model is based on LLaVA-Next architecture (Liu et al., 2024) with Qwen2-7B-Instruct (Yang et al., 2024) as the language model backbone and clip-vit-large-patch14-336 (Radford et al., 2021) as the vision encoder. The training consists of two stages. First, we pretrain the vision-language connector that aligns the outputs of clip-vit-large-patch14-336 to Qwen2-7B-Instruct, with the LLaVA LCS-558K [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain (Liu et al., 2023b; a). Then, we perform finetuning on PangeaIns, where we employ a learning rate of 2e-5, a batch size of 512, coupled with a cosine decay schedule with 0.03 warmup steps. We pretrain and finetune the model for 1 epoch, where pretraining took 4 hours with 8 H100 (32 GPU hours), and finetuning took 168 hours with 8 H100 (1344 GPU hours).
For evaluation, we compare Pangea against several state-of-the-art open source baselines, including English-centric models Llava-1.5-7B (Liu et al., 2023a), Llava-Next-7B (Liu et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), Cambrian-8B (Tong et al., 2024), Llava-OV-7B (Li et al., 2024b), Molmo-7B-D (Deitke et al., 2024) Llama3.2-11B (Dubey et al., 2024) and multilingual models PaliGemma-3B (Beyer et al., 2024), PALO-7B (Rasheed et al., 2025), mBLIP mT0-XL and mBLIP BLOOMZ (Geigle et al., 2024b). We also consider two text-only LLMs baselines Vicuna-1.5-7B (Zheng et al., 2023) and Qwen2-7B-Instruct (Yang et al., 2024), which are the backbones of Llava-Next and our Pangea respectively. We integrate our multimodal tasks in PangeaBench into lmms-eval (Li et al., 2024a), a multimodal evaluation package that supports many English multimodal benchmarks. We use lm-evaluation-harness (Biderman et al., 2024) to evaluate text-only tasks. We follow the original paper for their best models’ prompts in different tasks, and mostly reproduce their original numbers on datasets reported in the original papers.
@@ 4.1 Experimental Setup
$\Delta$ over SoTA Open +0.4 +10.9 -3.5 +7.3 -11.7 +18.3 -5.8 -4.2 +11.6 +12.3
@@
$\Delta$ over Best Open Model -0.2 +3.9 -0.1 +12.0 0.0 +9.4 -0.8 +2.3 +1.0 -3.7
@@ 5 Discussion
Role of English Data. In multilingual scenarios, English data plays a pivotal role in cross-lingual transfer. To investigate this, we sampled 500K examples from the translated data described in subsection 2.1, ensuring a consistent data distribution. We varied the ratio of English data while keeping the total number of training samples fixed at 500K. For the 17 multilingual languages in the translated subset, we evenly distributed the number of samples across languages.
@@
As shown in Figure 6 , English performance generally improves as the percentage of English data increases. More surprisingly, using no English data (full multilingual data) results in relatively lower multilingual performance. As we introduce more English data, multilingual performance improves, peaking at 38.7% with 40% English. However, performance drops sharply when English data reaches 100%. This suggests that English data aids cross-lingual transfer, however, over-reliance on it harms multilingual performance.
@@
Multimodal Chat in the Wild. One important application of MLLMs is to answer users’ queries in the wild. Here, we show the outputs of Pangea for the multimodal chat queries from our xChatBench. As shown in Appendix Figure 9, 10, 11, 12, 13, 14, Pangea successfully interprets the figures in different tasks and generates fluent and readable in certain languages. These qualitative examples further demonstrate the remarkable visual understanding ability of Pangea in multilingual contexts. On the other hand, we also identified a few bad cases shown in Figure 15, 16. Despite generating relevant responses to the queries, the model does not capture the key details of the images due to the lack of knowledge, which points out potential improvement directions in the future.
Preliminary Explorations of Multilingual OCR. Multilingual OCR emerged as a particularly challenging aspect of Pangea’s functionality. We made efforts to enhance its multilingual OCR capabilities. Further details on exploring multilingual OCR are available in Appendix J.